Show your solution completely
A system has constant volume (∆V=0) and the heat around the system increases by 50 J


A. What is the sign for heat (q) for the system?

B. What is ∆V equal to?

C. What is the value of internal energy of the system in Joules?


1
Expert's answer
2022-03-24T04:06:03-0400

Since the system has constant volume (ΔV=0) the term -PΔV=0 and work is equal to zero. Thus, in the equation ΔU=q+w w=0 and ΔU=q. The internal energy is equal to the heat of the system. The surrounding heat increases, so the heat of the system decreases because heat is not created nor destroyed. Therefore, heat is taken away from the system making it exothermic and negative. The value of Internal Energy will be the negative value of the heat absorbed by the surroundings.


a. negative (q<0)

b. ΔU=q + (-PΔV) = q+ 0 = q

c. ΔU = -45J